# Unisolver: PDE-Conditional Transformers Towards Universal Neural PDE Solvers

### v4

@@ 4.4 Model Analysis
Firstly, focusing on the effect of LLMs in Unisolver, we compared various LLMs and found comparable performances among them in Figure 8(a). Going further, we analyze the PDE embeddings in Figure 8(b-c), where principal component analysis (PCA) [14] is applied to intuitively visualize the LLM embeddings of equation symbols and deep PDE conditions learned by Unisolver for 1D time-dependent PDEs. In Figure 8(b), we observe that PDEs with similar complexity are encoded into similar embeddings, highlighting that LLM can indeed effectively capture prior mathematical knowledge. In Figure 8(c), the trajectories of deep conditions resemble parabolas with varying degrees of curvature, indicating that the learned deep conditions successfully capture the variations of PDE components.
@@ Appendix B Fine-tuning Performance
For 1D time-dependent PDEs, as shown in Figure 11, fine-tuning 100 epochs on the Burgers and Advection equations from PDEBench [37] significantly enhances Unisolver’s performance, reducing error by 61% compared to zero-shot results and achieving a 59.3% improvement over PDEformer [45] under the same fine-tuning conditions. These results prove the condition modeling in Unisolver is more effective than the computational graph proposed by PDEformer, especially for fast adaptation. For 2D mixed PDEs, as shown in Figure 12, after 200 epochs of fine-tuning for each dataset, Unisolver reduces error by more than 12% compared to zero-shot generalization performance and outperforms DPOT [11] under the same fine-tuning conditions by 14%, showcasing its ability to extract generalizable knowledge from diverse training datasets.
@@ Appendix E More Showcases
Figure 15: Error maps on the in-distribution test and zero-shot generalization (Burgers and Advection equation from PDEBench [37]) settings in 1D time-dependent PDEs. See Table 4 for numerical comparison. We visualize two cases: periodic boundary conditions and Robin boundary conditions in in-distribution tests. The number in the Burgers columns is the diffusion coefficient $\nu$ while the number in the Advection column is the advection speed $\beta$.
@@ Appendix F Analytical Solution for the String Vibration Equation
$u(x,t)=\frac{1}{2}(\underbrace{\Phi(x+\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}t)+\Phi(x-\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}t)}_{\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}Initial position}}})+\frac{1}{2\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}}\underbrace{\int^{x+\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}t}_{x-\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}t}\Psi(\xi)}_{\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}Initial velocity}}}\mathrm{d}\xi+\frac{1}{2\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}}\underbrace{\int_{0}^{t}\mathrm{d}\tau\int_{x-\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}(t-\tau)}^{x+\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}(t-\tau)}}_{\text{{\color[rgb]{0.45,0.45,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.45,0.45,0.9}Geometry}}}\underbrace{f(\xi,\tau)}_{\text{{\color[rgb]{0.45,0.45,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.45,0.45,0.9}Force}}}\mathrm{d}\xi,\$ (3)
where $\Phi(x)$, $\Psi(x)$ and $F(x,t)$ are odd, periodic functions with period $2{L}$ defined on the upper half plane, extended from $\phi(x)$, $\psi(x)$ and $f(x,t)$. The boundary conditions will be explicit by extending the equation to the upper half plane and solving it by operator splitting and characteristic lines.
Detailed proof can be found in [8] or other relevant books.
@@ Appendix G Benchmarks
We provide a detailed description of the three large-scale benchmarks in our experiments here: a challenging, self-generated heterogeneous 2D Navier-Stokes Equations dataset (HeterNS) and two large-scale benchmarks, one proposed by PDEformer [45], and the other collected by DPOT [11]. These benchmarks cover a wide range of PDEs and diverse generalization scenarios, which can test the generalizability of PDE solvers well.
@@ G.1 HeterNS
Similar to FNO [19], we consider the 2D Navier-Stokes equation in vorticity formulation for the viscous, incompressible fluid on a unit torus. We consider both in-distribution test and zero-shot generalization settings on HeterNS. See Figure 13 and 14 for a visual representation.
@@
This benchmark is proposed by PDEformer [45]. It contains 3 million high-quality 1D time-dependent PDEs with various equation components for training and then evaluates the model performance using in distribution test sets and zero-shot generalization performance on Burgers and Advection equation from PDEBench [37], which is another distinct benchmark. See Figure 15 for a visual representation.
@@
where $f_{i}(u)=c_{i1}u+c_{i2}u^{2}+c_{i3}u^{3},i=0,1$. Each coefficient $c_{ik}$ is set to zero with a probability of 0.5, and otherwise uniformly sampled from the interval $[-3,3]$. The variables $\kappa(x)$ and $s(x)$ can be zero, constant or physical fields, which are all randomly sampled from pre-defined distributions, as detailed in PDEformer’s original paper [45]. The initial condition $g(x)$ is randomly generated within the family of trigonometric functions, a super-position of sinusoidal waves as, $u_{0}(x)=\sum_{k_{i}=k_{1},...,k_{N}}A_{i}\sin(k_{i}x+\phi_{i})$, where $k_{i}=2\pi{n_{i}}/L_{x}$ are wave numbers and ${n_{i}}\in\mathbb{N}$ are selected randomly in $[1,n_{\mathrm{max}}]$, which is same as the zero-shot generalization tasks from PDEBench [37].
@@
Additionally, there is one important aspect to consider regarding the symbolic variations of equation symbols. A zero coefficient in the two polynomials $f_{i}$ results in the removal of a term from the equation. If the physical fields $\kappa(x)$ or $s(x)$ are zero, the corresponding term is removed from the prompt. When $\kappa(x)$ is constant, it is replaced by $\kappa$ to more accurately reflect the constant value, and the same applies to $s(x)$. These symbolic variations directly affect the equation formulations further embedded by the LLM, resulting in $5762^{6}\times 3\times 3=$ 576 types of LLM embeddings, corresponding to 576 distinct equation types.
@@
We employ the following two 1D PDE datasets from PDEBench [37] as zero-shot generalization tasks. All zero-shot generalization tasks follow periodic boundary conditions and the same initial condition family as the training dataset. The resolution of these samples is $1024\times 201$. For each dataset, we use 1000 test samples. We downsample the spatial resolution of these datasets to 256 and maintain the temporal resolution unchanged. The zero-shot PDEs consist of the Burgers equation and the Advection equation.
@@
Notably, the spatiotemporal domain of the equations in PDEBench is $[0,1]\times[0,2]$, whereas the training dataset uses the domain$[-1,1]\times[0,1]$. To directly infer from the model trained on 1D time-dependent PDEs, we need to align the spatiotemporal domains through spatial-temporal coordinate transformations, which will result in corresponding changes to the PDE components. Technically, the zero-shot PDEs after the coordinate transformation are given by:
@@
This benchmark is collected by DPOT [11], which consists of the following 12 diverse subsets from 4 benchmarks. We only conduct in-distribution tests in the 2D mixed PDEs. Notably, the in- distribution test set also involves challenging variations in the PDE components. See Figure 16 for a visual representations.
FNO-$\nu$ [19]
@@
PDEBench [37]
The following three subsets are derived from PDEBench [37], encompassing three distinct equations: the compressible Navier-Stokes equation (CNS), the diffusion-reaction equation (DR), and the shallow-water equation (SWE). All datasets considered in PDEBench adhere to periodic boundary conditions. The spatial resolution of this benchmark is $128\times 128$.
@@
In in-distribution test settings, flows are generated for each problem with different PDE components, which are a combination of three types: (1) boundary conditions, (2) fluid physical coefficients such as density and viscosity, and (3) the geometry of the field. The boundary conditions refer to the inlet velocity or movement velocity, depending on the specific case. Each type of PDE component corresponds to a distinct subset. In each subset, the corresponding PDE components are varied while other parameters remain constant. We mix the three subsets following DPOT’s configuration [11], resulting in 9,000 training samples and 1,000 testing samples. The initial resolution is $64\times 64$, which is then interpolated to $128\times 128$. The task is to predict the next 10 timesteps of velocity given the first 10 timesteps of observations.
@@ H.1 Loss and Metrics
Relative  $\text{Relative \ Promotion}=1-\frac{\epsilon_{\text{ours}}}{\epsilon_{\text{second-best model}}}.$ (14)
@@
Relative  $\text{Relative \ Drop}=\frac{\epsilon_{\text{ablation}}}{\epsilon_{\text{ours}}}-1.$ (15)
@@ H.2 Implementations for Each Benchmark
As outlined in Section 4, all the baseline models are trained under the same training strategy. We train the model using one-step predictions and test the model in an autoregressive manner. Specifically, all the models are trained for 300 epochs using the relative L2 loss and the ADAM optimizer [17] with an initial learning rate of $0.0005$, along with a cosine annealing learning rate scheduler [26]. The batch size is set to 60. After the training process, we use the checkpoint from the last epoch to evaluate the model performance.
We also provide the detailed model architecture hyperparameters in Table 12. We configure each model to align their model parameter numbers to ensure a fair comparison. Note that for MPP [2023], We utilize the parameter configuration of the tiny version containing approximately five million trainable parameters, which is comparable to Unisolver and other baselines.
The varying PDE components in this benchmark include the viscosity coefficient and external force. This physics information is provided to each baseline in an explicit or implicit way to ensure a fair comparison. For FNO [2021a], ViT [2020], FactFormer [2023b], and MPP [2023], we explicitly concatenate the viscosity coefficient and the external force to the model input along the channel dimension to ensure a fair comparison. As the viscosity coefficient is essentially a scalar, we repeat it along the spatial dimensions and then perform the channel-concatenating process. ICON [2023] is a special baseline which takes prompting trajectories as additional inputs to implicitly extract the physics information. Consequently, instead of providing the PDE components, we augment the input to ICON with five additional prompting trajectories with the same viscosity and external force as the target trajectory. Note that ICON also needs additional prompting trajectories when conducting evaluation. For PINO [2021b], we follow the experiment setting in the original paper and train the model with physics-informed loss as a soft regularization. The proportion of physics-informed loss with regard to data loss is set to 0.1.
@@
The model we use to compare with PDEformer-L contains 19M trainable parameters, which is comparable to the 22M parameters of PDEformer-L. The model scalability experiments in Section 4.4 also show model configurations with different number of trainable parameters. We progressively increase the Unisolver parameter from 3M to 63M, resulting in 4 different model configurations. We present the detailed configurations of these models in Table 14. Note that in this benchmark, we utilize an adapted version PolyINR [35] to decode the encoder output from the Transformer backbone.
@@ H.3 Details of the LLM embeddings
Note that the pre-training dataset of PDEformer [45] contains the PDE family following the formulation:
@@ I.4 Full Scalability
As a supplement to Figure 10 in the main text, we also conduct experiments on different zero-shot generalization tasks from [37] and record the concrete data in Table 21 for clarity.
@@ I.5 Efficiency Analysis
We provide the inference time and memory consumption for each model to predict a single frame on the HeterNS, along with the calculation time and memory consumption of the numerical solver, which is a pseudo-spectral solver adopted by FNO [19] used to generate the HeterNS dataset, to calculate the next frame, as summarized in the Table 22. The results are measured on an A100 GPU with a batch size of 1. Unisolver demonstrates comparable inference speed to FNO, while consuming less memory. Besides, all neural PDE solvers are approximately 1,000 times faster than the numerical solver, highlighting their potential as efficient surrogate models.

### v5

@@ 4.4 Model Analysis
Firstly, focusing on the effect of LLMs in Unisolver, we compared various LLMs and found comparable performances among them in Figure 8(a). Going further, we analyze the PDE embeddings in Figure 8(b-c), where principal component analysis (PCA) (Jolliffe & Cadima, 2016) is applied to intuitively visualize the LLM embeddings of equation symbols and deep PDE conditions learned by Unisolver for 1D time-dependent PDEs. In Figure 8(b), we observe that PDEs with similar complexity are encoded into similar embeddings, highlighting that LLM can indeed effectively capture prior mathematical knowledge. In Figure 8(c), the trajectories of deep conditions resemble parabolas with varying degrees of curvature, indicating that the learned deep conditions successfully capture the variations of PDE components.
@@ Appendix B Fine-tuning Performance
For 1D time-dependent PDEs, as shown in Figure 11, fine-tuning 100 epochs on the Burgers and Advection equations from PDEBench (Takamoto et al., 2022) significantly enhances Unisolver’s performance, reducing error by 61% compared to zero-shot results and achieving a 59.3% improvement over PDEformer (Ye et al., 2024) under the same fine-tuning conditions. These results prove the condition modeling in Unisolver is more effective than the computational graph proposed by PDEformer, especially for fast adaptation. For 2D mixed PDEs, as shown in Figure 12, after 200 epochs of fine-tuning for each dataset, Unisolver reduces error by more than 12% compared to zero-shot generalization performance and outperforms DPOT (Hao et al., 2024) under the same fine-tuning conditions by 14%, showcasing its ability to extract generalizable knowledge from diverse training datasets.
@@ Appendix E More Showcases
Figure 15: Error maps on the in-distribution test and zero-shot generalization (Burgers and Advection equation from PDEBench (Takamoto et al., 2022)) settings in 1D time-dependent PDEs. See Table 4 for numerical comparison. We visualize two cases: periodic boundary conditions and Robin boundary conditions in in-distribution tests. The number in the Burgers columns is the diffusion coefficient $\nu$ while the number in the Advection column is the advection speed $\beta$.
@@ Appendix F Analytical Solution for the String Vibration Equation
$u(x,t)=\frac{1}{2}(\underbrace{\Phi(x+\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}t)+\Phi(x-\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}t)}_{\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}Initial position}}})+\frac{1}{2\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}}\underbrace{\int^{x+\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}t}_{x-\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}t}\Psi(\xi)}_{\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}Initial velocity}}}\mathrm{d}\xi+\frac{1}{2\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}}\underbrace{\int_{0}^{t}\mathrm{d}\tau\int_{x-\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}(t-\tau)}^{x+\text{{\color[rgb]{0.37,0.67,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.37,0.67,1}a}}(t-\tau)}}_{\text{{\color[rgb]{0.45,0.45,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.45,0.45,0.9}Geometry}}}\underbrace{f(\xi,\tau)}_{\text{{\color[rgb]{0.45,0.45,0.9}\definecolor[named]{pgfstrokecolor}{rgb}{0.45,0.45,0.9}Force}}}\mathrm{d}\xi,\\$ (3)
where $\Phi(x)$, $\Psi(x)$ and $F(x,t)$ are odd, periodic functions with period 2L2{L} defined on the upper half plane, extended from $\phi(x)$, $\psi(x)$ and $f(x,t)$. The boundary conditions will be explicit by extending the equation to the upper half plane and solving it by operator splitting and characteristic lines.
Detailed proof can be found in (Evans, 2022) or other relevant books.
@@ Appendix G Benchmarks
We provide a detailed description of the three large-scale benchmarks in our experiments here: a challenging, self-generated heterogeneous 2D Navier-Stokes Equations dataset (HeterNS) and two large-scale benchmarks, one proposed by PDEformer (Ye et al., 2024), and the other collected by DPOT (Hao et al., 2024). These benchmarks cover a wide range of PDEs and diverse generalization scenarios, which can test the generalizability of PDE solvers well.
@@ G.1 HeterNS
Similar to FNO (Li et al., 2021a), we consider the 2D Navier-Stokes equation in vorticity formulation for the viscous, incompressible fluid on a unit torus. We consider both in-distribution test and zero-shot generalization settings on HeterNS. See Figure 13 and 14 for a visual representation.
@@
This benchmark is proposed by PDEformer (Ye et al., 2024). It contains 3 million high-quality 1D time-dependent PDEs with various equation components for training and then evaluates the model performance using in distribution test sets and zero-shot generalization performance on Burgers and Advection equation from PDEBench (Takamoto et al., 2022), which is another distinct benchmark. See Figure 15 for a visual representation.
@@
where $f_{i}(u)=c_{i1}u+c_{i2}u^{2}+c_{i3}u^{3},i=0,1$. Each coefficient $c_{ik}$ is set to zero with a probability of 0.5, and otherwise uniformly sampled from the interval [3,3][-3,3]. The variables $\kappa(x)$ and $s(x)$ can be zero, constant or physical fields, which are all randomly sampled from pre-defined distributions, as detailed in PDEformer’s original paper (Ye et al., 2024). The initial condition $g(x)$ is randomly generated within the family of trigonometric functions, a super-position of sinusoidal waves as, $u_{0}(x)=\sum_{k_{i}=k_{1},...,k_{N}}A_{i}\sin(k_{i}x+\phi_{i})$, where $k_{i}=2\pi{n_{i}}/L_{x}$ are wave numbers and $n_{i}$$\in\mathbb{N}$ are selected randomly in $1,n_{\mathrm{max}}]$, which is same as the zero-shot generalization tasks from PDEBench (Takamoto et al., 2022).
@@
Additionally, there is one important aspect to consider regarding the symbolic variations of equation symbols. A zero coefficient in the two polynomials $f_{i}$ results in the removal of a term from the equation. If the physical fields $\kappa(x)$ or $s(x)$ are zero, the corresponding term is removed from the prompt. When $\kappa(x)$ is constant, it is replaced by $\kappa$ to more accurately reflect the constant value, and the same applies to $s(x)$. These symbolic variations directly affect the equation formulations further embedded by the LLM, resulting in $2^{6}\times 3\times 3=576$ types of LLM embeddings, corresponding to 576 distinct equation types.
@@
We employ the following two 1D PDE datasets from PDEBench (Takamoto et al., 2022) as zero-shot generalization tasks. All zero-shot generalization tasks follow periodic boundary conditions and the same initial condition family as the training dataset. The resolution of these samples is $1024\times 201$. For each dataset, we use 1000 test samples. We downsample the spatial resolution of these datasets to 256 and maintain the temporal resolution unchanged. The zero-shot PDEs consist of the Burgers equation and the Advection equation.
@@
[…]
@@
This benchmark is collected by DPOT (Hao et al., 2024), which consists of the following 12 diverse subsets from 4 benchmarks. We only conduct in-distribution tests in the 2D mixed PDEs. Notably, the in- distribution test set also involves challenging variations in the PDE components. See Figure 16 for a visual representations.
FNO-$\nu$ (Li et al., 2021a)
@@
PDEBench (Takamoto et al., 2022)
The following three subsets are derived from PDEBench (Takamoto et al., 2022), encompassing three distinct equations: the compressible Navier-Stokes equation (CNS), the diffusion-reaction equation (DR), and the shallow-water equation (SWE). All datasets considered in PDEBench adhere to periodic boundary conditions. The spatial resolution of this benchmark is $128\times 128$.
@@
In in-distribution test settings, flows are generated for each problem with different PDE components, which are a combination of three types: (1) boundary conditions, (2) fluid physical coefficients such as density and viscosity, and (3) the geometry of the field. The boundary conditions refer to the inlet velocity or movement velocity, depending on the specific case. Each type of PDE component corresponds to a distinct subset. In each subset, the corresponding PDE components are varied while other parameters remain constant. We mix the three subsets following DPOT’s configuration (Hao et al., 2024), resulting in 9,000 training samples and 1,000 testing samples. The initial resolution is $64\times 64$, which is then interpolated to $128\times 128$. The task is to predict the next 10 timesteps of velocity given the first 10 timesteps of observations.
@@ H.1 Loss and Metrics
$\text{Relative \ Promotion}=1-\frac{\epsilon_{\text{ours}}}{\epsilon_{\text{second-best model}}}.$ (14)
@@
$\text{Relative \ Drop}=\frac{\epsilon_{\text{ablation}}}{\epsilon_{\text{ours}}}-1.$ (15)
@@ H.2 Implementations for Each Benchmark
As outlined in Section 4, all the baseline models are trained under the same training strategy. We train the model using one-step predictions and test the model in an autoregressive manner. Specifically, all the models are trained for 300 epochs using the relative L2 loss and the ADAM optimizer (Kingma & Ba, 2015) with an initial learning rate of $0.0005$, along with a cosine annealing learning rate scheduler (Loshchilov & Hutter, 2016). The batch size is set to 60. After the training process, we use the checkpoint from the last epoch to evaluate the model performance.
We also provide the detailed model architecture hyperparameters in Table 12. We configure each model to align their model parameter numbers to ensure a fair comparison. Note that for MPP (2023), We utilize the parameter configuration of the tiny version containing approximately five million trainable parameters, which is comparable to Unisolver and other baselines.
The varying PDE components in this benchmark include the viscosity coefficient and external force. This physics information is provided to each baseline in an explicit or implicit way to ensure a fair comparison. For FNO (2021a), ViT (2020), FactFormer (2023b), and MPP (2023), we explicitly concatenate the viscosity coefficient and the external force to the model input along the channel dimension to ensure a fair comparison. As the viscosity coefficient is essentially a scalar, we repeat it along the spatial dimensions and then perform the channel-concatenating process. ICON (2023) is a special baseline which takes prompting trajectories as additional inputs to implicitly extract the physics information. Consequently, instead of providing the PDE components, we augment the input to ICON with five additional prompting trajectories with the same viscosity and external force as the target trajectory. Note that ICON also needs additional prompting trajectories when conducting evaluation. For PINO (2021b), we follow the experiment setting in the original paper and train the model with physics-informed loss as a soft regularization. The proportion of physics-informed loss with regard to data loss is set to 0.1.
@@
The model we use to compare with PDEformer-L contains 19M trainable parameters, which is comparable to the 22M parameters of PDEformer-L. The model scalability experiments in Section 4.4 also show model configurations with different number of trainable parameters. We progressively increase the Unisolver parameter from 3M to 63M, resulting in 4 different model configurations. We present the detailed configurations of these models in Table 14. Note that in this benchmark, we utilize an adapted version PolyINR (Singh et al., 2023) to decode the encoder output from the Transformer backbone.
@@ H.3 Details of the LLM embeddings
Note that the pre-training dataset of PDEformer (Ye et al., 2024) contains the PDE family following the formulation:
@@ I.4 Full Scalability
As a supplement to Figure 10 in the main text, we also conduct experiments on different zero-shot generalization tasks from (Takamoto et al., 2022) and record the concrete data in Table 21 for clarity.
@@ I.5 Efficiency Analysis
We provide the inference time and memory consumption for each model to predict a single frame on the HeterNS, along with the calculation time and memory consumption of the numerical solver, which is a pseudo-spectral solver adopted by FNO (Li et al., 2021a) used to generate the HeterNS dataset, to calculate the next frame, as summarized in the Table 22. The results are measured on an A100 GPU with a batch size of 1. Unisolver demonstrates comparable inference speed to FNO, while consuming less memory. Besides, all neural PDE solvers are approximately 1,000 times faster than the numerical solver, highlighting their potential as efficient surrogate models.